OpenStudy (kittiwitti1):

Ok, I don't know if I should put this in the calc1 subforums but LaTex is below

OpenStudy (kittiwitti1):

\({\lim_{x \rightarrow 0}}\frac{1-\cos{h}}{h}\) at \(\large h=\pm0.1,\pm0.01,\pm0.001\)

OpenStudy (sweetburger):

this is basically asking \[\lim_{x \rightarrow 0^+}\frac{ 1-\cos(h) }{ h }\] just plugin the points and test them.

OpenStudy (agent0smith):

^Yep, since the function is defined at all those values, just plug them in and evaluate. They want you to be able to see that the original limit approaches zero.

OpenStudy (kittiwitti1):

Oh, this managed to get bumped? lol it didn't show up on mine -- Sorry, I was saving this for another problem

zepdrix (zepdrix):

Subforums? Do people still use those? :3

OpenStudy (kittiwitti1):

Okay so \[\lim_{x\rightarrow-3^{-}}\frac{2e^{x}}{(x+3)^{3}}\] they want the infinite limit which I calculated via graph to be \(-\infty\) but the answer that the professor gave was "as \(x\rightarrow3^{-}...\)" \(2e^{x}\rightarrow2e^{-3}\) (positive #s) \((x+3)^{3}\rightarrow0^{-}\) (negative #s)

OpenStudy (kittiwitti1):

the limit is \(x\rightarrow-3^{-}\) btw

OpenStudy (agent0smith):

Use \Large to make it more legible\[\Large \lim_{x\rightarrow-3^{-}}\frac{2e^{x}}{(x+3)^{3}}\]

OpenStudy (kittiwitti1):

oh yeah sorry;; so idk, did I get it at least half right with \(f(x)\rightarrow-\infty\)?

OpenStudy (agent0smith):

Yes, it's negative infinity, since the numerator is positive and denominator is negative and approaching zero.

OpenStudy (kittiwitti1):

I honestly thought it was more of an end behavior question idk. But then there's this -- and the answer to another question \(\Large{\lim_{x\rightarrow\frac{\pi}{2}^{+}}\sec{x}}\)

OpenStudy (agent0smith):

End behaviour means when x is approaching pos. or neg. infinity.

OpenStudy (kittiwitti1):

So end behavior \(\ne\) infinite limits?

OpenStudy (agent0smith):

End behaviour refers to when x is approaching pos. or neg. infinity. Nothing to do with a limit that is equal to a value approaching infinity.

OpenStudy (kittiwitti1):

... okay, then what does the WA link mean by \(-\infty\)?

OpenStudy (kittiwitti1):

It's all so confusing sorry ;;

OpenStudy (agent0smith):

That the function value approaches neg. infinity, when x approaches -3 from left.

OpenStudy (kittiwitti1):

So end behavior is the x-value and infinite limit is the y-value or output?

OpenStudy (kittiwitti1):

*x-value/input

OpenStudy (agent0smith):

End behaviour will look like\[\Large \lim_{x \rightarrow \infty}\]or\[\Large \lim_{x \rightarrow -\infty} \]

OpenStudy (agent0smith):

And yes.

OpenStudy (kittiwitti1):

OH. Well that was simpler than I thought... wow

OpenStudy (kittiwitti1):

I'm still not sure why the problem with e has a two-part answer though

OpenStudy (kittiwitti1):

\(\color{blue}{\text{Originally Posted by}}\) kittiwitti1 \[\lim_{x\rightarrow-3^{-}}\frac{2e^{x}}{(x+3)^{3}}\] they want the infinite limit which I calculated via graph to be \(-\infty\) but the answer that the professor gave was "as \(x\rightarrow3^{-}...\)" \(2e^{x}\rightarrow2e^{-3}\) (positive #s) \((x+3)^{3}\rightarrow0^{-}\) (negative #s) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (agent0smith):

He's just describing the result.

OpenStudy (kittiwitti1):

Oh, so it's still \(f(x)\rightarrow-\infty\) then?

OpenStudy (agent0smith):

\[\Large \lim_{x\rightarrow-3^{-}}\frac{2e^{x}}{(x+3)^{3}} = -\infty\]and then what he wrote is just describing why

OpenStudy (kittiwitti1):

http://prntscr.com/cd78wi Oh, alright. Thank you for clarifying

OpenStudy (kittiwitti1):

If anyone sees this, btw... how do I calculate asymptotes with a function like \(\Large{\frac{2e^{x}}{(x+3)^{3}}}\)?

zepdrix (zepdrix):

In the land of math, dividing by 0 is a crime.\[\large\rm \frac{1}{0}\qquad \leftarrow\text{this is not a number}\] Asymptotic behavior leads to this though. \[\large\rm \frac{1}{x}\qquad\text{asymptote at x=0}\] \[\large\rm \frac{1}{x-2}\qquad \text{asymptote at x=2}\]

zepdrix (zepdrix):

To state simply, the denominator tells us about asymptotes. Whatever value causes the denominator to be 0 leads to an asymptote.

zepdrix (zepdrix):

So in your problem, \(\large\rm (x+3)^3=0\) Solving for x will tell us that special number that causes the denominator to be 0. Or in this case, simply eye-balling it, because -3 and 3 give us 0, ya? :)

OpenStudy (kittiwitti1):

@zepdrix I was actually wondering about the numerator in terms of finding any possible horizontal or slant asymptotes, but that was also helpful :p

zepdrix (zepdrix):

I don't think you can get slant/oblique asymptotes from non-polynomial stuff. Those types of asymptotes show up when the degree of the numerator is exactly one larger than the degree of the denominator. Some examples of graphs with oblique asymptotes:\[\large\rm \frac{x^4+4}{x^3+7},\qquad\qquad \frac{x^2}{x+1}\] So you don't need to worry about that when exponential stuff is thrown into the mix.

zepdrix (zepdrix):

Horizontal asymptotes show up at the far ends of the graph, end behavior as you were talking about before :)

OpenStudy (kittiwitti1):

Soooo ... when the numerator has a wacky \(x\) then disregard all HA/SA rules? hahah

OpenStudy (kittiwitti1):

I mean... I WA'd it and got this http://www.wolframalpha.com/input/?i=(2e%5Ex)%2F%5B(x%2B3)%5E3%5D+asymptotes

zepdrix (zepdrix):

Disregard SA rules, yes. But take for example something with wacky x stuff like this,\[\large\rm \frac{e^{-7x}}{x^2+8}\] As x approaches the end of the graph (on the right side), e^{-7x} is approaching zero much much faster than x^2+8 is blowing up. So the numerator is winning toward 0, so horizontal asymptote at y=0. I'm not sure what techniques you've covered... It would some fancier tools than algebra to determine horizontal asymptotes for exponential stuff.

zepdrix (zepdrix):

I guess I didn't need to come up with another example. Yours would have worked just fine. It's a comparison, a race, between the numerator and denominator. You want to know who is winning the race.

zepdrix (zepdrix):

When you plug x=-3bajillion into the numerator and denominator separately, \(\large\rm 2e^{-3bajillion}=\tiny \text{really really small}\) \(\large\rm (-3bajillion+3)^3=\text{ehh sort of large}\) Numerator is winning to zero.

zepdrix (zepdrix):

Ok that's bad explanation maybe XD haha.

OpenStudy (kittiwitti1):

\(\Large{bajillion}\) LOL

zepdrix (zepdrix):

Ok to summarize: Vertical Asymptotes: Set denominator equal to 0. Horizontal Asymptotes: Look at the ends of the graph. Slant Asymptotes: None if wacky stuff included.

OpenStudy (kittiwitti1):

(mom's insulting me with rather crude things and telling me this site is "the reason you get C D F" ok) Can you calculate HA from the fcn alone?

zepdrix (zepdrix):

lolol

zepdrix (zepdrix):

fcn?

OpenStudy (kittiwitti1):

Function. Abbreviated lol

zepdrix (zepdrix):

You can get a clear idea of what the HA should be by plugging in a really really big x value (positive or negative). But you'll have to use this idea of limits to be completely satisfied.

OpenStudy (kittiwitti1):

; - ;

OpenStudy (kittiwitti1):

I just used the normal n < m rule for HA = 0 in this function though\[\lim_{x\rightarrow8^{+}}\frac{1}{x-8}\]

zepdrix (zepdrix):

Ok fair enough :) You could also plug in a really big x value to get an idea of what's going on off in the distance. If we plug x=7batrillion... or no no, let's use a more sensible number like x=10007. So we get,\[\large\rm \frac{1}{x-8}\quad=\quad \frac{1}{10007-8}=\frac{1}{9999}\approx.0001\]It's getting really close to 0 when x is big.

OpenStudy (kittiwitti1):

Hm, makes sense

zepdrix (zepdrix):

meow -_-

OpenStudy (kittiwitti1):

Okay so... the n < m cannot apply to the function that has numerator with exponential \(x\)?

zepdrix (zepdrix):

Correct.\[\large\rm \frac{x^n+3}{x^m+94}\]Where n<m, collapses to 0 for big x. \[\large\rm \frac{(x^n+3)e^{5x}\cos(7x)}{x^m+94}\]But this is a whole something else going on here D:

OpenStudy (kittiwitti1):

what the heck is that a failed experiment x'D

zepdrix (zepdrix):

ya haha :D

OpenStudy (kittiwitti1):

ok well I'm glad I didn't have that lol

OpenStudy (agent0smith):

Cool beanios.